# Left-left-right-right magnetic order in spin-1/2 Kitaev-Heisenberg chain

### v1

@@ Abstract
In this work, we perform a combination of analytical and numerical studies on the phase diagram of the spin-1/2 Kitaev-Heisenberg chain in the region of negative Kitaev and positive Heisenberg couplings. Apart from the antiferromagnetic phase, we find a magnetically ordered phase with left-left-right-right order and a gapless phase with central charge value $c=1$, resolving the existing contradictory results in literature regarding this parameter region. In particular, the origin of the left-left-right-right phase is clarified based on a perturbative Luttinger liquid analysis. Using a coupled-chain method, we further show that the one-dimensional (1D) left-left-right-right order is able to give a quasi-1D explanation for the 2D stripy order of the same model on the honeycomb lattice.
@@ I Introduction
One-dimensional (1D) generalized Kitaev spin models can be constructed from their two-dimensional (2D) counterparts Kitaev2006 ; Nayak2008 ; Jackeli2009 ; Chaloupka2010 ; Rau2014 ; Kimchi2014 ; Wang2017 by selecting one row out of the honeycomb lattice, where the word “generalized” is used to emphasize the fact that additional interactions beyond the Kitaev interaction are also included in the model in order to describe real Kitaev materials Witczak-Krempa2014 ; Rau2016 ; Winter2017 ; Hermanns2018 . Recently, motivated by the potential of the 1D studies in providing hints for the 2D Kitaev physics, there have been surging research efforts in studying 1D generalized Kitaev models Sela2014 ; Gruenewald2017 ; Agrapidis2018 ; Agrapidis2019 ; Catuneanu2019 ; You2020 ; You2020b ; Yang2019 ; Yang2020 ; Yang2020b ; Yang2021b ; Luo2021 ; Luo2021b ; Sorensen2021 ; Yang2022a ; Yang2022_2 ; Yang2022 ; Yang2022d ; Yang2022e ; Yang2022f ; Zhao2022 ; Laurell2022 . On the other hand, in addition to being helpful for understanding 2D physics, 1D generalized Kitaev models contain intriguing strongly correlated physics on their own, and it has been established that these 1D models have complicated nonsymmorphic symmetry group structures Yang2022a ; Yang2022e , leading to exotic properties including emergent SU(2)1 conformal invariance Yang2019 ; Yang2022d ; Yang2022e , nonsymmorphic bosonization Yang2022_2 ; Yang2022 , nonlocal string order parameters Catuneanu2019 ; Luo2021 , and magnetic phases breaking nonsymmorphic symmetries Yang2020 .
@@
Motivated by the above questions, in this work, we perform a combination of analytical and numerical studies in the region $\phi\in(\arctan{2},\pi/2)$, in which $K=-\sin(\phi)$ and $J=\cos(\phi)$ are the Kitaev and Heisenberg interactions, respectively. Notice that the hidden SU(2) symmetric FM point is located at $\phi=\arctan(2)$. By a combination of perturbative Luttinger liquid analysis and density matrix renormalization group (DMRG) numerical simulations, we find that there is a magnetically ordered phase close to the $K=-2J$ point shown by “LLRR” in the phase diagram in Fig. 1 (a), where the spin ordering exhibits a collinear left-left-right-right (LLRR) pattern, which is consistent with the finding in Ref. Agrapidis2018, (named as staggered-$xy$ order there) and different from the spiral phase identified in Ref. Yang2020, . The key spirit of the strategy is straightforward to understand. After a unitary transformation called four-sublattice rotation, the Hamiltonian can be decomposed into two parts, one part the easy-plane XXZ model, and the other part a four-site periodic term. Since the Luttinger parameter diverges when the system approaches the $K=-2J$ point, the system is very sensitive to perturbations, and as a result, driven into an ordered state by the four-site periodic term. The predictions are supported by our DMRG White1992 ; White1993 ; Schollwock2011 simulations.
Furthermore, our DMRG numerics indicate that there exists a phase transition point $\phi_{c1}$ which separates the LLRR phase from a $c=1$ gapless phase. Therefore, unlike the claims in the literatures Sela2014 ; Agrapidis2018 ; Yang2020 , the system have both ordered and gapless phases in the region $\phi\in(\arctan(2),\pi/2)$. Since the $c=1$ phase is close to the FM Kitaev point which has exponentially large ground state degeneracy, the numerics in the gapless phase are very difficult. An analytical understanding of the gapless phase remains unclear and will be left for future investigations.
We make two extensions of the LLRR phase in the 1D spin-1/2 Kitaev-Heisenberg model. First, by weakly coupling an infinite number of chains on the honeycomb lattice, we recover the stripy phase in the 2D Kitaev-Heisenberg Chaloupka2013 ; Rau2014 , thereby providing a quasi-1D understanding to the 2D stripy phase. Second, by adding a Gamma term, we find that the LLRR phase extends to a finite region in the phase diagram of the spin-1/2 Kitaev-Heisenberg-Gamma chain shown in Fig. 1 (b).
The rest of the paper is organized as follows. In Sec. II, the model Hamiltonian is introduced and the symmetry group is analyzed. In Sec. III, the magnetic order in the 1D spin-1/2 Kitaev-Heisenberg model is determined by a combination of perturbative Luttinger liquid analysis and DMRG numerics. Sec. IV generalizes the analysis to the Kitaev-Heisenberg-Gamma chain. In Sec. V, the FM phase in the FM Heisenberg region is discussed based on a Luttinger liquid analysis. Sec. VI provides DMRG numerical evidence to the analytical results in the preceding sections. Sec. VII is devoted to a derivation of the stripy phase in the 2D Kitaev-Heisenberg model using a coupled chain method. Sec. VIII summarizes the main results of the paper.
@@
Figure 1: Schematic plot of the phase diagrams of the 1D spin-1/2 (a) Kitaev-Heisenberg model and (b) Kitaev-Heisenberg-Gamma model in the parameter region $\phi\in(0,\pi)$. In both (a,b), in terms of the parametrization in Eq. (2), the $J$, FM, $-K$, $J$ points on the horizontal axis have $\phi$-values 00, $\arctan(2)$, $\pi/2$, $\pi$, respectively, where FM is the hidden SU(2) symmetric point satisfying $K=-2J<0$. In (b), the vertical axis is $\Gamma$, and because of the equivalence in Eq. (6), only the $\Gamma>0$ region is shown.
@@ II.1 Model Hamiltonian
$H_{0}=\sum_{<ij>\in\gamma\,\text{bond}}\big{[}KS_{i}^{\gamma}S_{j}^{\gamma}+J\vec{S}_{i}\cdot\vec{S}_{j}\big{]},$ (1)
@@
in which “Sublattice $i$” ($1\leq i\leq 4$) represents all the sites $i+4n$ ($n\in\mathbb{Z}$) in the chain, and the spin operators $S^{\alpha}$ ($\alpha=x,y,z$) are abbreviated as $\alpha$ for simplicity. It can be verified that the Hamiltonian $H^{\prime}_{0}=U_{4}H_{0}U_{4}^{-1}$ in the four-sublattice rotated frame is of the same form as Eq. (1) except that $K$ and $J$ should be replaced by $K+2J$, $-J$, respectively. Hence $U_{4}$ establishes the following equivalence relation in the parameter space,
@@
Particularly, $U_{4}$ reveals two hidden SU(2) symmetric points $K=-2J<0$ and $K=-2J>0$, corresponding to FM and antiferromagnetic (AFM) Heisenberg model, respectively. From here on, we will call the frames before and after the $U_{4}$ transformation as the original and $U_{4}$ frames, respectively.
@@ II.2 Symmetries
in which $T$ is time reversal; $T_{na}$ ($n\in\mathbb{Z}$) is translation by $n$ lattice sites; $I$ is the spatial inversion with the inversion center located at site 00; and $R(\hat{n},\theta)$ denotes the global spin rotation around the $\hat{n}$-direction by an angle $\theta$. The symmetry group $G_{0}$ of the Kitaev-Heisenberg chain is
@@
Since $T_{4a}=[R(\hat{z},-\frac{\pi}{2})T_{a}]^{4}$ is a symmetry element and the abelian group $\langle T_{4a}\rangle$ is a normal subgroup of $G_{0}$, it is legitimate to consider the quotient group $G_{0}/\langle T_{4a}\rangle$.  [^1] One can actually consider $G_{0}/\langle T_{2a}\rangle$ for the Kitaev-Heisenberg model, but $T_{4a}$ is considered instead so that we can conveniently compare $G_{0}/\langle T_{4a}\rangle$ with the symmetry group of the Kitaev-Heisenberg-Gamma model in the $U_{4}$ frame later. It has been proved in Ref. Yang2020, that $G_{0}/\langle T_{4a}\rangle$ is isomorphic to $(\mathbb{Z}_{2}\times\mathbb{Z}_{2})\ltimes D_{4h}$, in which $\mathbb{Z}_{2}\times\mathbb{Z}_{2}=\langle T_{a}I,T_{2a}\rangle/\langle T_{4a}\rangle$ and $D_{4h}=\langle T\rangle\times D_{4}$, where $D_{4}$ is the dihedral group of order $8$ given by $D_{4}=\langle R(\hat{z},-\frac{\pi}{2})T_{a},R(\hat{y},\pi)T_{a}I\rangle/\langle T_{4a}\rangle$. We note that an intuitive way to understand the origin of the $D_{4}$ group is by observing that if the spatial operations $T_{a}$ and $T_{a}I$ are removed from $R(\hat{z},-\frac{\pi}{2})T_{a}$ and $R(\hat{y},\pi)T_{a}I$, then the actions in the spin space $\langle R(\hat{z},-\frac{\pi}{2}),R(\hat{y},\pi)\rangle$ generate the symmetry group of a spin square shown in Fig. 3, which is exactly the $D_{4}$ group. As a consequence of the above analysis of the group structure of $G_{0}$, we have the short exact sequence
@@
The group $G_{0}$ is nonsymmorphic in the sense that the above short exact sequence is non-splitting, i.e., there is no way to embed the quotient group $G_{0}/\langle T_{4a}\rangle$ into $G_{0}$.
@@
in which $\hat{n}_{1}=\frac{1}{\sqrt{2}}(1,-1,0)^{T}$ and $R(\hat{n}_{1},\pi)T_{a}=R(\hat{y},\pi)R(\hat{z},-\frac{\pi}{2})T_{a}$. The group $G_{1}$ satisfies $G_{1}/\langle T_{4a}\rangle=\mathbb{Z}_{2}\ltimes C_{4h}$, in which $\mathbb{Z}_{2}=\langle T_{a}I\rangle$ and $C_{4h}=\langle T\rangle\times C_{4}$ where $C_{4}=\langle R(\hat{n}_{1},\pi)T_{a}\rangle/\langle T_{4a}\rangle$ is the cyclic group of order $4$. Therefore, there is the short exact sequence
@@
We note that $G_{1}$ and $G_{1}^{\prime}$ is isomorphic to each other since they are the symmetry groups of the same model albeit in different frames. This is indeed true, as can be seen from the isomorphism $\mathbb{Z}_{2}\ltimes C_{4h}\cong D_{4h}$.
For all above nonsymmorphic symmetry groups, the non-splitting property can be most easily seen by considering the subgroup $\langle R(\hat{z},-\frac{\pi}{2})T_{a}\rangle$ and the following short exact sequence,
@@ II.3 Phase diagram
Before going to technical discussions, we give a description of the phase diagram shown in Fig. 1. We first consider the phase diagram of the Kitaev-Heisenberg model shown in Fig. 1 (a). The AFM Heisenberg point $J$, FM Heisenberg point $-J$, and the FM Kitaev point $-K$ are located at $\phi=0$, $\phi=\pi$, and $\phi=\pi/2$, respectively. There is another special point FM at $\phi=\arctan(2)$, which is a hidden SU(2) symmetric point satisfying $K=-2J<0$. We note that the region $\phi\in(\arctan(2),\pi/2)$ is related to $\phi\in(\pi/2,\pi)$ by the $U_{4}$ transformation, hence the physics in these two regions are unitarily equivalent.
The phase diagram in Fig. 1 (a) divides into five phases: AFM phase between $J$ and FM; LLRR phase between FM and $\phi_{c1}\approx 0.4\pi$; a gapless $c=1$ phase between $\phi_{c1}$ and $-K$, where $c$ represents the central charge; another $c=1$ phase between $-K$ and $\phi_{c2}\approx 0.56\pi$, which is dual to the $c=1$ phase to the left of the $-K$ point in Fig. 1 (a) under the $U_{4}$ transformation; and an FM phase between $\phi_{c2}$ and $-J$, which is dual to the LLRR phase in the region $\phi\in[\text{FM}^{\prime},\phi_{c1}]$ under the $U_{4}$ transformation. The spin ordering has a four-site periodicity in the LLRR phase, aligning along $\hat{x}$, $\hat{x}$, $-\hat{x}$, $-\hat{x}$ or $\hat{y}$, $\hat{y}$, $-\hat{y}$, $-\hat{y}$ directions in a four-site unit cell, hence the name LLRR (left-left-right-right for short). We note that the LLRR phase was erroneously identified as a spiral phase in Ref. Yang2020, . The numerics become very difficult in the $c=1$ phase, since the system is very close to the FM Kitaev point, which is exactly solvable and has an exponentially large infinite ground state degeneracy in the thermodynamic limit (degeneracy is $2^{L/2}$ for a system of $L$ sites). Whether the low energy theory of the $c=1$ phase is a Luttinger liquid with an emergent U(1) symmetry or two Majorana fermions without emergent U(1) symmetry remains unclear, which will be left for future investigations.
By adding a small Gamma interaction, nearly all the phases extend to a finite area in the two-parameter phase diagram as shown in Fig. 1 (b). In particular, the LLRR phase on the $\Gamma=0$ line extends to the d-LLRR phase, where “d-LLRR” is “distorted-left-left-right-right” for short, the name of which comes from the fact that the spins are no longer collinear in this case. Whether the gapless $c=1$ phase on the horizontal axis extends to a finite region remains unclear, which is denoted as “KSL” (“Kitaev spin liquid” for short) in Fig. 1 (b) and will be left for future studies. When $\Gamma$ becomes larger, there appears a Luttinger liquid phase in the positive $J$ region denoted as “LL” in Fig. 1, which has been discussed in detail in Ref. Yang2020, .

### v2

@@ Abstract
In this work, we perform a combination of analytical and numerical studies on the phase diagram of the spin-1/2 Kitaev-Heisenberg chain in the region of negative Kitaev and positive Heisenberg couplings. Apart from the antiferromagnetic phase, we find a magnetically ordered phase with left-left-right-right order and a gapless phase with central charge value $c=1$, resolving the existing contradictory results in literature regarding this parameter region. In particular, the origin of the left-left-right-right order is clarified based on a perturbative Luttinger liquid analysis. The left-left-right-right phase is further shown to persist in the Kitaev-Heisenberg-Gamma model when a small nonzero Gamma interaction is introduced. Using a coupled-chain method, we also demonstrate that the one-dimensional (1D) left-left-right-right order gives a quasi-1D explanation for the 2D stripy order of the same model on the honeycomb lattice.
@@ I Introduction
One-dimensional (1D) generalized Kitaev spin models can be constructed from their two-dimensional (2D) counterparts Kitaev2006 ; Nayak2008 ; Jackeli2009 ; Chaloupka2010 ; Rau2014 ; Kimchi2014 ; Wang2017 by selecting one row out of the honeycomb lattice, where the word “generalized” is used to emphasize the fact that additional interactions beyond the Kitaev interaction are also included in the model in order to describe real Kitaev materials Witczak-Krempa2014 ; Rau2016 ; Winter2017 ; Hermanns2018 . Recently, motivated by the potential of the 1D studies in providing hints for the 2D Kitaev physics, there have been surging research efforts in studying 1D generalized Kitaev models Sela2014 ; Gruenewald2017 ; Agrapidis2018 ; Agrapidis2019 ; Catuneanu2019 ; You2020 ; You2020b ; Yang2019 ; Yang2020 ; Yang2020b ; Yang2021b ; Luo2021 ; Luo2021b ; Sorensen2023 ; Sorensen2021 ; Yang2022a ; Yang2022_2 ; Yang2022 ; Yang2022d ; Yang2022e ; Yang2022f ; Zhao2022 ; Laurell2022 . On the other hand, in addition to being helpful for understanding 2D physics, 1D generalized Kitaev models contain intriguing strongly correlated physics on their own, and it has been established that these 1D models have complicated nonsymmorphic symmetry group structures Yang2022a , leading to exotic properties including emergent SU(2)1 conformal invariance Yang2019 ; Yang2022d ; Yang2022e , nonsymmorphic bosonization Yang2022_2 ; Yang2022 , nonlocal string order parameters Catuneanu2019 ; Luo2021 , solitons Sorensen2023 , and magnetic phases breaking nonsymmorphic symmetries Yang2020 .
@@
Figure 1: Schematic plot of the phase diagram of the 1D spin-1/2 Kitaev-Heisenberg-Gamma model, where the horizontal axis (i.e., the $\Gamma=0$ line) corresponds to the Kitaev-Heisenberg model. The phase diagram in the region $\phi\in(0,\pi)$ close to the $\Gamma=0$ line is shown, where the coordinate $\phi$ of the horizontal axis is defined in Eq. (2) as $K=-\sin(\phi)$, $J=\cos(\phi)$. The $J$, FM, $-K$, $J$ points on the horizontal axis have $\phi$-values 00, $\arctan(2)$, $\pi/2$, $\pi$, respectively, where FM is the hidden SU(2) symmetric point satisfying $K=-2J<0$. The vertical axis is $\Gamma$, and because of the equivalence in Eq. (6), only the $\Gamma>0$ region is shown. In the $\Gamma\neq 0$ region, “d-LLRR” is “distorted-left-left-right-right” for short; “LL” is “Luttinger liquid” for short; and “KSL” is “Kitaev spin liquid” for short, in which region the nature of the physics remains to be explored.
Motivated by the above questions, in this work, we perform a combination of analytical and numerical studies in the region $\phi\in(\arctan{2},\pi/2)$, in which $K=-\sin(\phi)$ and $J=\cos(\phi)$ are the Kitaev and Heisenberg interactions, respectively. Notice that the hidden SU(2) symmetric FM point is located at $\phi=\arctan(2)$. By a combination of perturbative Luttinger liquid analysis and density matrix renormalization group (DMRG) numerical simulations, we find that there is a magnetically ordered phase close to the $K=-2J$ point shown by “LLRR” on the horizontal axis of the phase diagram in Fig. 1, where the spin ordering exhibits a collinear left-left-right-right (LLRR) pattern, which is consistent with the finding in Ref. Agrapidis2018, (named as staggered-$xy$ order there) and different from the spiral phase identified in Ref. Yang2020, . The key spirit of the strategy is straightforward to understand. After a unitary transformation called four-sublattice rotation, the Hamiltonian can be decomposed into two parts, one part the easy-plane XXZ model, and the other part a four-site periodic term. Since the Luttinger parameter diverges when the system approaches the $K=-2J$ point, the system is very sensitive to perturbations, and as a result, driven into an ordered state by the four-site periodic term. The predictions are supported by our DMRG White1992 ; White1993 ; Schollwock2011 simulations.
Furthermore, as shown in Fig. 2, our DMRG numerics indicate that there exists a phase transition point $\phi_{c1}\approx 0.4\pi$ which separates the LLRR phase from a $c=1$ gapless phase. Therefore, unlike the claims in the literatures Sela2014 ; Agrapidis2018 ; Yang2020 , the system have both ordered and gapless phases in the region $\phi\in(\arctan(2),\pi/2)$. Since the $c=1$ phase is close to the FM Kitaev point which has exponentially large ground state degeneracy, the numerics in the gapless phase are very difficult. An analytical understanding of the gapless phase remains unclear and will be left for future investigations.
Figure 2: Central charge values of the spin-1/2 Kitaev-Heisenberg chain as a function of $\phi$. DMRG numerics are performed for $L=48,100,256$ sites using open boundary conditions. The bond dimension $m$ and truncation error $\epsilon$ in DMRG calculations are taken as $m=1000$, $\epsilon=10^{-9}$.
We make two extensions of the LLRR phase in the 1D spin-1/2 Kitaev-Heisenberg model. First, by adding a small nonzero Gamma term, we find that the LLRR phase extends to a finite region in the phase diagram of the spin-1/2 Kitaev-Heisenberg-Gamma chain, having a non-collinear magnetic ordering pattern, which can be viewed as a slightly distorted version of the LLRR ordering, denoted as “d-LLRR” in Fig. 1, where “d” is “distorted” for short. Second, by weakly coupling an infinite number of chains on the honeycomb lattice, we recover the stripy phase in the 2D Kitaev-Heisenberg Chaloupka2013 ; Rau2014 , thereby providing a quasi-1D understanding to the 2D stripy phase.
The rest of the paper is organized as follows. In Sec. II, the model Hamiltonian is introduced and the symmetry group is analyzed. In Sec. III, the LLRR and $c=1$ phases in the 1D spin-1/2 Kitaev-Heisenberg model are studied by perturbative Luttinger liquid analysis and DMRG numerics. Sec. IV generalizes the analysis to the Kitaev-Heisenberg-Gamma chain, and shows that the LLRR phase extends to a finite region in the phase diagram of the Kitaev-Heisenberg-Gamma model when a nonzero Gamma term is introduced. Sec. V briefly discusses phases in the phase diagram of the 1D Kitaev-Heisenberg-Gamma model other than the LLRR phase, especially the derivation of the FM phase using a perturbative Luttinger liquid analysis. Sec. VI is devoted to a derivation of the stripy phase in the 2D Kitaev-Heisenberg model using a coupled chain method based on the 1D results. In Sec. VII, we summarize the main results of the paper.
@@ II.1 Model Hamiltonian
$H_{0}=\sum_{<ij>\in\gamma\,\text{bond}}\big{(}KS_{i}^{\gamma}S_{j}^{\gamma}+J\vec{S}_{i}\cdot\vec{S}_{j}\big{)},$ (1)
@@
in which “Sublattice $i$” ($1\leq i\leq 4$) represents all the sites $i+4n$ ($n\in\mathbb{Z}$) in the chain, and the spin operators $S^{\alpha}$ ($\alpha=x,y,z$) are abbreviated as $\alpha$ for simplicity. It can be verified that the Hamiltonian $H^{\prime}_{0}=U_{4}H_{0}U_{4}^{-1}$ in the four-sublattice rotated frame is of the same form as Eq. (1), except that $K$ and $J$ should be replaced by $K+2J$ and $-J$, respectively. Hence $U_{4}$ establishes the following equivalence relation in the parameter space,
@@
Particularly, $U_{4}$ reveals two hidden SU(2) symmetric points $K=-2J<0$ and $K=-2J>0$, corresponding to FM and antiferromagnetic (AFM) Heisenberg models, respectively. From here on, we will call the spin coordinate systems before and after the $U_{4}$ transformation as the original and $U_{4}$ frames, respectively.
@@ II.2.1 Kitaev-Heisenberg model
in which $T$ is the time reversal operation; $T_{a}$ is the translation operator by one site, and $T_{na}$ is translation by $n$ lattice sites where $n$ is an integer; $I$ is the spatial inversion with the inversion center located at site 00; and $R(\hat{n},\theta)$ denotes the global spin rotation around the $\hat{n}$-direction by an angle $\theta$. The symmetry group $G_{0}$ of the Kitaev-Heisenberg chain is
@@
Since $T_{4a}=[R(\hat{z},-\frac{\pi}{2})T_{a}]^{4}$ is a symmetry element and the abelian group $\langle T_{4a}\rangle$ is a normal subgroup of $G_{0}$, it is legitimate to consider the quotient group $G_{0}/\langle T_{4a}\rangle$. Here we note that one can actually consider $G_{0}/\langle T_{2a}\rangle$ for the Kitaev-Heisenberg model, but $T_{4a}$ is considered instead so that we can conveniently compare $G_{0}/\langle T_{4a}\rangle$ with the symmetry group of the Kitaev-Heisenberg-Gamma model in the $U_{4}$ frame to be discussed later.
It has been proved in Ref. Yang2020, that $G_{0}/\langle T_{4a}\rangle$ is isomorphic to $(\mathbb{Z}_{2}\times\mathbb{Z}_{2})\ltimes D_{4h}$, in which $\mathbb{Z}_{2}\times\mathbb{Z}_{2}=\langle T_{a}I,T_{2a}\rangle/\langle T_{4a}\rangle$ and $D_{4h}=\langle T\rangle\times D_{4}$, where $D_{4}$ is the dihedral group of order $8$ given by $D_{4}=\langle R(\hat{z},-\frac{\pi}{2})T_{a},R(\hat{y},\pi)T_{a}I\rangle/\langle T_{4a}\rangle$. We note that an intuitive way to understand the origin of the $D_{4}$ group is by observing that if the spatial operations $T_{a}$ and $T_{a}I$ are removed from $R(\hat{z},-\frac{\pi}{2})T_{a}$ and $R(\hat{y},\pi)T_{a}I$, then the group $\langle R(\hat{z},-\frac{\pi}{2}),R(\hat{y},\pi)\rangle$ generated by actions in the spin space represents the symmetry group of a spin square shown in Fig. 4, which is exactly the $D_{4}$ group. As a consequence of the above analysis of the group structure of $G_{0}$, we have the short exact sequence
@@
Particularly, the group $G_{0}$ is nonsymmorphic in the sense that the short exact sequence in Eq. (10) is non-splitting, i.e., there is no way to embed the quotient group $G_{0}/\langle T_{4a}\rangle$ into $G_{0}$.
@@ II.2.2 Kitaev-Heisenberg-Gamma model
in which $\hat{n}_{1}=\frac{1}{\sqrt{2}}(1,-1,0)^{T}$. The group $G_{1}$ satisfies $G_{1}/\langle T_{4a}\rangle=\mathbb{Z}_{2}\ltimes C_{4h}$, in which $\mathbb{Z}_{2}=\langle T_{a}I\rangle$ and $C_{4h}=\langle T\rangle\times C_{4}$ where $C_{4}=\langle R(\hat{n}_{1},\pi)T_{a}\rangle/\langle T_{4a}\rangle$ is the cyclic group of order $4$. Therefore, there is the short exact sequence
@@
It is straightforward to see that $G_{1}^{\prime}$ is a subgroup of $G_{0}$ by comparing Eq. (14) with Eq. (10), where $D_{4h}$ is a subgroup of $(\mathbb{Z}_{2}\times\mathbb{Z}_{2})\ltimes D_{4h}$. We note that $G_{1}$ and $G_{1}^{\prime}$ are isomorphic to each other since they are the symmetry groups of the same model albeit in different frames. This is indeed true, as can be seen from the isomorphism $\mathbb{Z}_{2}\ltimes C_{4h}\cong D_{4h}$.
II.2.3 Nonsymmorphic nature of the symmetry groups
For all above nonsymmorphic symmetry groups in Sec. II.2 and Sec. II.2.2, the non-splitting property can be most easily seen by considering the subgroup $\langle R(\hat{z},-\frac{\pi}{2})T_{a}\rangle$ and the following short exact sequence,
@@ II.3 Phase diagram
Before going to technical discussions, we give a description of the phase diagram shown in Fig. 1.
II.3.1 Kitaev-Heisenberg model
We first describe the phase diagram of the Kitaev-Heisenberg model corresponding to the horizontal axis in Fig. 1. The AFM Heisenberg point $J$, FM Heisenberg point $-J$, and the FM Kitaev point $-K$ are located at $\phi=0$, $\phi=\pi$, and $\phi=\pi/2$, respectively. There is another special point FM at $\phi=\arctan(2)$, which is a hidden SU(2) symmetric point satisfying $K=-2J<0$. We note that the region $\phi\in(\arctan(2),\pi/2)$ is related to $\phi\in(\pi/2,\pi)$ by the $U_{4}$ transformation, hence the physics in these two regions are unitarily equivalent.
The horizontal axis in Fig. 1 divides into five phases: AFM phase between $J$ and FM; LLRR phase between FM and $\phi_{c1}\approx 0.4\pi$; a gapless $c=1$ phase between $\phi_{c1}$ and $-K$, where $c$ represents the central charge; another $c=1$ phase between $-K$ and $\phi_{c2}\approx 0.56\pi$, which is dual to the $c=1$ phase to the left of the $-K$ point in Fig. 1 under the $U_{4}$ transformation; and an FM phase between $\phi_{c2}$ and $-J$, which is dual to the LLRR phase in the region $\phi\in[\text{FM}^{\prime},\phi_{c1}]$ under the $U_{4}$ transformation.
The spin ordering has a four-site periodicity in the LLRR phase in the original frame, aligning along $\hat{x}$, $\hat{x}$, $-\hat{x}$, $-\hat{x}$ or $\hat{y}$, $\hat{y}$, $-\hat{y}$, $-\hat{y}$ directions in a four-site unit cell, hence the name LLRR (left-left-right-right for short). We note that the LLRR phase was erroneously identified as a spiral phase in Ref. Yang2020, . The numerics become very difficult in the $c=1$ phase, since the system is very close to the exactly solvable FM Kitaev point, which has an exponentially large infinite ground state degeneracy in the thermodynamic limit (degeneracy is $2^{L/2}$ for a system of $L$ sites) Brzezicki2007 ; You2008 . More detailed analytical and numerical analysis of the $c=1$ phase will be left for future investigations.
II.3.2 Kitaev-Heisenberg-Gamma model
By adding a small Gamma interaction, nearly all the phases extend to a finite area in the two-parameter phase diagram as shown in Fig. 1. In particular, the LLRR phase on the $\Gamma=0$ line extends to the d-LLRR phase, where “d-LLRR” is “distorted-left-left-right-right” for short, the name of which comes from the fact that the pattern of the spin ordering can be viewed as a slightly distorted version of the LLRR ordering in Kitaev-Heisenberg model, no longer collinear when the Gamma term is nonzero. On the other hand, whether the gapless $c=1$ phase on the horizontal axis extends to a finite region remains unclear, which is denoted as the “KSL” (“Kitaev spin liquid” for short) phase in Fig. 1 and will be left for future studies. When $\Gamma$ becomes larger, there appears a Luttinger liquid phase in the positive $J$ region denoted as “LL” in Fig. 1, which has been discussed in detail in Ref. Yang2020, .